OpenStudy (anonymous):

how would i solve this? The sum of 3 consecutive integral numbers is 117. Find the numbers.

OpenStudy (anonymous):

x+(x+1)+(x+2)=117

OpenStudy (anonymous):

x+(x+1)+(x+2)=117

OpenStudy (anonymous):

n+(n+1)+(n+2)=117 (where n is the smallest integer) n+n+1+n+2=117 3n+3=117 3n=114 n= 38 the numbers are 38, 39, and 40

OpenStudy (anonymous):

n+(n+1)+(n+2)=117 so n= 38